Identify the greater number, wherever possible, in each of the following?
2^10 or 10^2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The expansion of 210= 2 × 2 × 2 × 2 × 2 × 2 × 2 × 2 × 2 × 2 = 1024
The expansion of 10^2= 10 × 10= 100
Clearly,
1024 > 100
So, 2^10> 10^2
Hence 28is the greater number.

Thanks